OpenStudy (anonymous):

State the coordinates of the center and the length of the radius given the equation of the circle below. x^2+(y+4)^2=3

OpenStudy (anonymous):

First define the equation of a circle

OpenStudy (anonymous):

\[(x+h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

notice is says "state" rather than "find" because it is in exactly the form you want it in:\[(x-h)^2+(y-k)^2=r^2\] match them up

OpenStudy (anonymous):

You should be able to get it easily now :)

OpenStudy (anonymous):

Thankyou guys ! :)

OpenStudy (anonymous):

You're welcome